OpenStudy (anonymous):

What is the recursive formula of A(n)=2+9(n-1) A. A(n)=A(n-1)+2 B. A(n)=A(n-1)+9 C. A(n)=A(n-1)-9 D. A(n)=A(n-1)-2

OpenStudy (amistre64):

write out a few terms; it looks arithmetical

OpenStudy (amistre64):

since we are tacking on the 9 times tables; id say its the +9 route

OpenStudy (anonymous):

I was thinking that but , wasnt sure

OpenStudy (amistre64):

a0=2 a1 = 2+9 a2 = (2+9)+9 a3 = (2+9+9)+9 a4 = (2+9+9+9)+9 hmm,

OpenStudy (anonymous):

ok,thanks

OpenStudy (amistre64):

youre welcome, and when in doubt, write it out. brute math is useful at times :)